Use the order of operations to find 12 when you have 8,5,3,1? please help me ive been trying to figure this out forever
Answer by stanbon(75887)

Use the order of operations to find 12 when you have 8,5,3,1? please help me ive been trying to figure this out forever
------
(8-5)(3+1) = 3*4 = 12
===========================
Cheers,
Stan H.